### v1

# Bidirectional predictive coding

Abstract
Predictive coding (PC) is an influential computational model of visual learning and inference in the brain. Classical PC was proposed as a top-down generative model, where the brain actively predicts upcoming visual inputs, and inference minimises the prediction errors. Recent studies have also shown that PC can be formulated as a discriminative model, where sensory inputs predict neural activities in a feedforward manner. However, experimental evidence suggests that the brain employs both generative and discriminative inference, while unidirectional PC models show degraded performance in tasks requiring bidirectional processing. In this work, we propose bidirectional PC (bPC), a PC model that incorporates both generative and discriminative inference while maintaining a biologically plausible circuit implementation. We show that bPC matches or outperforms unidirectional models in their specialised generative or discriminative tasks, by developing an energy landscape that simultaneously suits both tasks. We also demonstrate bPC’s superior performance in two biologically relevant tasks including multimodal learning and inference with missing information, suggesting that bPC resembles biological visual inference more closely.
1 Introduction
Visual inference plays a critical role in the brain, providing information processing for interpreting and interacting with the environment. Two frameworks have emerged to explain how the brain could implement visual inference. The first describes vision as a bottom-up discriminative process, which filters sensory inputs to extract behaviorally relevant information, akin to inference in feedforward neural networks [1]. The other formalises vision as a generative process where the brain generates top-down predictions of the sensory inputs, and develops neural responses as a result of Bayesian inference [2]. Experimental evidence suggests that visual perception may be a combination of both frameworks [3, 4]. However, most computational models for visual inference are either discriminative or generative. Models that are a combination of both, often have to compromise on either supervised or unsupervised learning performance, both of which are hallmarks of visual learning. Hence, it is unknown how the brain is able to perform these diverse tasks accurately.
In this work, we propose bidirectional predictive coding (bPC), a novel model of biological vision that can perform both generative and discriminative inference based on predictive coding. Traditional predictive coding can be formulated as either generative or discriminative, and in both cases it matches biological data and the performance of analogous machine learning models. It also employs Hebbian rules to learn the model parameters supporting either generative or discriminative inference, making it biologically more plausible than models relying on backpropagation (BP) to learn. In this work, we focus on evaluating and understanding bPC’s performance in a range of computational tasks, where predictive coding models with either generative or discriminative inference have been individually successful. Our key contributions can be summarised as follows:
  • We propose bidirectional predictive coding, a biologically plausible model of visual perception employing both generative and discriminative inference that naturally arises from minimising a single energy function;
  • We show that bPC performs as well as its purely discriminative or generative counterparts on both supervised classification tasks and unsupervised representation learning tasks, and outperforms precedent hybrid predictive coding models;
  • We provide an explanation for the superior performance of bPC in both tasks, by comparing it to unidirectional models and showing that it learns an energy landscape which benefits detecting out-of-distribution data points;
  • We further show that bPC outperforms other PC models in two biologically relevant tasks, including multimodal learning and inference with occluded visual scenes, indicating its potential as a more faithful model for visual inference in the brain.
2 Background and related work
Figure 1: Model architectures used in our experiments. Arrows indicate the direction of prediction. Dashed arrows represent initialisation. Discriminative models (red) perform bottom-up predictions, generative models (blue) perform top-down predictions, and bidirectional models incorporate both. Models in the top row implement local predictive coding computations, whereas those in the bottom row rely on non-local, BP-based computations. Each prediction arrow corresponds to bidirectional connections in a neural circuit: one for prediction, the other for error backpropagation.
Discriminative predictive coding.
Recent studies have focused on PC models with bottom-up predictions from sensory inputs to latent activities [5, 6], similar to feedforward neural networks. The energy function of this type of PC models is given by:
$E_{disc}(x,V)=\sum\nolimits_{l=2}^{L}\frac{1}{2}\|x_{l}-V_{l-1}f(x_{l-1})\|_{2}^{2}$ (1)
In this energy function, $x_{l}$ is the neural activity of layer $l$ where $x_{1}$ is set to the sensory data and $x_{L}$ is set to a label. The parameter $V_{l}$ represents the bottom-up weights of layer $l$ and $f$ is an activation function. We refer to this model as discriminative predictive coding (discPC) and illustrate it in the top panel of Figure 1A. In discPC, inference is performed by a forward pass from $x_{1}$ to $x_{l}$’s, and the learning rule of this model approximates BP, using only local computations and Hebbian plasticity [5, 7]. Recent work showed that discPC performs comparably to BP in classifying MNIST, Fashion-MNIST and CIFAR-10 [8] and outperforms BP at learning scenarios encountered by the brain, such as online and continual learning [6]. However, discPC lacks unsupervised learning capabilities due to the non-uniqueness of the solution to its generative dynamics [9].
Generative predictive coding.
Classically, predictive coding (PC) uses top-down predictions from neural activities to sensory data, based on the hypothesis that the brain learns by minimizing the error between its predicted sensory inputs and the actual sensory input [10, 11]. With neurons arranged hierarchically, the objective function, or energy [6], of PC can be written as:
$E_{gen}(x,W)=\sum\nolimits_{l=1}^{L-1}\frac{1}{2}\|x_{l}-W_{l+1}f(x_{l+1})\|_{2}^{2}$ (2)
where $x_{1}$ is set to sensory data and $x_{L}$ can be clamped to a label for supervised learning or remain free for unsupervised learning. The parameters $W_{l}$ are top-down weights. We refer to this formulation of predictive coding as generative predictive coding (genPC). In genPC, inference of latent activities $x_{l}$ is performed by clamping $x_{1}$ to sensory inputs and iteratively updating $x_{l}$’s via gradient descent on the energy $E_{gen}$. genPC’s top-down predictions are illustrated in the top-left panel of Figure 1B.
GenPC has explained various visual phenomena, such as extra-classical receptive field effects and repetition suppression [10, 12, 13, 14]. More recently, genPC has been employed to model associative memory [15] and unsupervised image generation [16, 17, 8]. The computational framework of genPC can be implemented within neural networks that utilize local computations and Hebbian plasticity [18]. However, genPC has poor performance in supervised learning tasks, where $x_{1}$ is clamped to the input image and the corresponding label is iteratively inferred in $x_{L}$ [19].
Hybrid predictive coding.
In this work, we benchmark our bPC model primarily against hybrid predictive coding (hybridPC). hybridPC is the only plausible PC model that incorporates both slow, iterative inference and fast, feedforward inference, to date. In HybridPC, fast inference is enabled by a bottom-up network added to the genPC network architecture, which serves only to initialize neural activities and does not affect their dynamics, as illustrated in the top-right panel of Figure 1B. The neural dynamics and weight updates of hybridPC minimise the following energy function:
$E_{hybrid}(x,W,V)=\sum\nolimits_{l=1}^{L-1}\frac{1}{2}\|x_{l}-W_{l+1}f(x_{l+1})\|_{2}^{2}+\sum\nolimits_{l=2}^{L}\frac{1}{2}\|\textrm{sg}(x_{l})-V_{l-1}f(\textrm{sg}(x_{l-1}))\|_{2}^{2}$ (3)
where sg($\cdot$) is the stop gradient operation. HybridPC can perform both supervised and unsupervised learning. However, the supervised learning performance of hybridPC is much poorer than that of discPC. This is due to the stop gradient operation in the bottom-up stream of this model that prevents it from learning a discriminative model that maps well from $x_{1}$ to $x_{L}$ [19].
Other predictive coding models.
Other formulations of PC models combine discriminative and generative characteristics. Sun and Orchard [9] noted the energy-minimising nature of PC models could theoretically allow image generation in discPC (Eq. 1), by clamping $x_{L}$ to a label and iteratively updating $x_{1}$. However, the generated images appear nonsensical due to the non-unique inferential dynamics solutions. Even with decay mechanisms to make the solution unique, the output generations remain highly noisy [9]. Noting that any neuron in a PC model can be updated to minimise an energy function regardless of the model architecture, Salvatori et al. [20] generalized the idea of Sun and Orchard [9] to a PC model where all neurons are interconnected with each other. However, it only slightly outperforms the classification performance of a linear classifier. Both studies suggest that dedicated circuits for discriminative and generative tasks are needed for satisfactory performance. Finally, Qiu et al. [21] proposed a PC model that performs feedforward, bottom-up inference using the transpose of the top-down weights. Although this model performs well in both supervised and unsupervised tasks, it is unlikely that the brain shares synaptic connections between distinct streams of processing, making it an implausible model compared to bPC.
Discriminative and generative models of the brain.
Computational models of cortical processing can be roughly categorized into discriminative, such as those that use a feedforward neural network to filter visual inputs [22, 23], and generative, which are rooted in the Bayesian Brain hypothesis [24, 25] and instantiated by theories like predictive coding [10, 11]. However, increasing evidence suggests that the brain takes a combined approach [26, 3, 4]. For example, the brain might use a discriminative approach to rapidly extract important features from visual scenes, and perform Bayesian inference on these features to form high-level perceptions such as object categories [27]. Only a small number of models explicitly integrates these two approaches. The Helmholtz machine, an early model trained with the wake-sleep algorithm [28], is analogous to a discriminative-generative process and can be seen as a precursor to variational autoencoders [29] trained via BP. The Symmetric Predictive Estimator [30] is a recent model designed to capture bidirectional processing in the brain. However, its effectiveness beyond simple toy tasks remains unclear.
3 Model: bidirectional predictive coding
In contrast to genPC and discPC, bPC neurons perform both top-down and bottom-up predictions, as shown in the top panel of Figure 1C. Using the notations previously introduced, the energy function of bPC is given by:
$E(x,W,V)=\sum\nolimits_{l=1}^{L-1}\frac{\alpha_{gen}}{2}\|x_{l}-W_{l+1}f(x_{l+1})\|_{2}^{2}+\sum\nolimits_{l=2}^{L}\frac{\alpha_{disc}}{2}\|x_{l}-V_{l-1}f(x_{l-1})\|_{2}^{2}$ (4)
where $W_{l}$ are the top-down weights, $V_{l}$ are the bottom-up weights, and $f$ is an activation function. $\alpha_{gen}$ and $\alpha_{disc}$ are scalar weighting constants, which are needed to account for magnitude differences in the errors of bottom-up and top-down predictions. The weighting constants can be viewed as learnable precision parameters [11]; however, they are tuned and kept constant in our implementation for simplicity. Unlike in hybridPC, no stop gradient is applied. In each trial of learning, the neural activities are first updated to minimise $E$ via gradient descent (neural dynamics) following:
$\frac{dx_{l}}{dt}=-\nabla_{x}E=-\epsilon_{l}^{gen}-\epsilon_{l}^{disc}+f^{\prime}(x_{l})\odot\Big{(}W_{l}^{\top}\epsilon_{l-1}^{gen}+V^{\top}_{l}\epsilon_{l+1}^{disc}\Big{)},$ (5)
where
$\epsilon_{l}^{gen}:=\alpha_{gen}(x_{l}-W_{l+1}f(x_{l+1})),\quad\epsilon_{l}^{disc}:=\alpha_{disc}(x_{l}-V_{l-1}f(x_{l-1}))$ (6)
denote the top-down and bottom-up prediction errors of neurons in layer $l$ respectively. $f^{\prime}$ denotes the derivative of the function $f$, and $\odot$ is the element-wise product.
After updating neural activities, the weights are updated to minimise $E$ via a gradient descent step:
$\Delta W_{l}\propto-\nabla_{W_{l}}E=\epsilon_{l-1}^{gen}f(x_{l})^{\top},\quad\Delta V_{l}\propto-\nabla_{V_{l}}E=\epsilon_{l+1}^{disc}f(x_{l})^{\top}$ (7)
Similar to discPC [5] and hybridPC [19], we initialise the model layers using a feedforward sweep from layer $x_{1}$ to $x_{L}$ along the bottom-up predictions. For instance, the second and third layers are initialised as $x_{2}=V_{1}f(x_{1})$ and $x_{3}=V_{2}f(V_{1}f(x_{1}))$ respectively. This initialisation strategy can be interpreted as a mechanism for fast amortized inference when a sensory input is initially encountered, which is also observed in the brain [31, 26].
3.1 Neural implementation
Figure 2: Neural implementation of bPC (left) and discPC (right).
The computations described in Eqs 5 and 7 can be implemented in a neural network with fully local computation and plasticity, as illustrated in Figure 2. This network contains value neurons, which encode $x_{l}$, error neurons, which represent prediction errors, and synaptic connections, which encode the model parameters. All computations are local, relying solely on pre- and post- synaptic activity. Value neuron dynamics depend on local error signals, their own activity, and incoming synaptic weights. Similarly, error neuron activity depends only on adjacent value neurons and synaptic weights. The plasticity of the weights is also Hebbian, as it is a product of pre- and post-synaptic activity. The local implementation of bPC inherits that of genPC or discPC (shown in Figure 2 right), although it has two distinct error neurons per value neuron for bottom-up and top-down prediction errors.
3.2 Flexible learning
bPC can be trained both in supervised and unsupervised settings. In all cases, neurons in intermediate layers (layer $2$ to $L-1$) are un-clamped and evolve according to the neural dynamics given in equation 5. In supervised settings, the first layer $x_{1}$ is clamped to the input, while the top layer $x_{L}$ is clamped to the target labels. In unsupervised settings, $x_{L}$ is left un-clamped, allowing bPC to learn compressed representations of the input. A mixed setting is also possible, where only a subset of neurons in $x_{L}$ are clamped to label information, while others remain free. In this setting, the model can jointly infer labels and learn an associated compressed representation.
4 Experiments
4.1 Supervised classification and generation
Figure 3: bPC accurately classifies and generates class average images on MNIST and Fashion-MNIST. A: Training set-up. The models are trained with $x_{1}$ fixed to images and $x_{L}$ fixed to labels. B: Classification accuracy of models. C: Examples of the generated images conditional on class labels (left) and RMSEs between generated images and mean images of each class (right).
Firstly, we demonstrate that a single bPC model simultaneously performs discriminative and generative tasks. In this experiment, bPC learns to associate images and labels, such that it can classify an input image and generate an average image of a class when a label is provided. We compare bPC with discPC, genPC, hybridPC, and their backpropagation equivalents on MNIST and Fashion-MNIST datasets for the same architecture with two hidden layers (256 neurons each). Additional PC models are evaluated in the supplementary material, SM. During training, we clamped layer $x_{1}$ to images and layer $x_{L}=x_{4}$ to their corresponding one-hot labels (Figure 3A). Post-training, we evaluated discriminative performance by measuring classification accuracy given only the image. Specifically, we fixed layer $x_{1}$ to the image and inferred labels at layer $x_{L}$ after 100 inference steps. Generative performance was assessed by the root mean squared error (RMSE) between the image generated from each class label and the corresponding average image for that class. We obtained generated images in $x_{1}$ after fixing $x_{L}$ to a class label and performing a 100-step iterative inference. A detailed training and evaluation procedure for all experiments, including hyperparameter selection, is provided in SM.
A comparison of classification accuracy and generative RMSE across models is shown in Figure 3B. bPC achieves a classification accuracy comparable to discPC and discBP for both datasets. Conversely, genBP and hybridPC show lower accuracy. Regarding generative performance, bPC exhibits RMSE scores similar to genPC, hybridPC, and genBP, while discPC demonstrates significantly higher generation errors. Generated images from discPC contain minimal class-relevant features, whereas bPC and other models produce clear, representative images for each class (Figure 3C).
These results confirm previously reported results that unidirectional PC models excel only in their specialized domain: discPC fails in generation due to non-unique learning minima [9], while genPC and hybridPC have a poor classification accuracy [19]. In contrast, bPC performs well in generative and discriminative tasks simultaneously. Notably, bPC uses only half the neurons needed to maintain two pathways for generative and discriminative tasks separately, which can be more energy efficient.
We also found that discriminative weighting ($\alpha_{disc}$) needs to be set higher than the generative weighting due to the larger magnitude of top-down prediction errors. The exploration into the impact of $\alpha_{disc}/\alpha_{gen}$ is presented in SM.
4.2 Unsupervised representation learning
Figure 4: bPC outperforms other PC models at learning low-dimensional representations. A: Training set-up, where only $x_{1}$ is clamped to input images. B: Reconstruction RMSE across models and datasets. C: Example reconstructions.
Next, we show that bPC learns compressed representations in the absence of supervision. We evaluated the representation learning performance of bPC, genPC, hybridPC, and their BP equivalents across tasks of increasing complexity, using MNIST, Fashion-MNIST, CIFAR-10, and CIFAR-100 datasets. discPC was not evaluated due to its inability to perform unsupervised learning. We trained models where only layer $x_{1}$ was clamped to input images, leaving all other layers free (see Figure 4A). MNIST and Fashion-MNIST models had two hidden layers (256 neurons each) and a representation layer $x_{L}$ of 30 neurons. Training was constrained to 8 inference steps between weight updates to assess the models’ capacity for fast inference. CIFAR models had five convolutional layers and a representation layer of 256 neurons, trained with 32 inference steps. Activity decay was used at layer $x_{L}$ to stabilize learning and constrain representations. After training, representations were obtained in $x_{L}$ after 100 inference steps for a given input image in $x_{1}$. Images were reconstructed using 100 inference steps after clamping $x_{L}$ to a representation and reinitialising layers $x_{1}$ to $x_{L-1}$. We measured the RMSE between the original images and their reconstructions from representations.
Figure 4B shows that bPC consistently outperforms genPC by achieving a lower RMSE. On MNIST and Fashion-MNIST, bPC matches the performance of hybridPC. On more complex datasets such as CIFAR-10 and CIFAR-100, bPC significantly surpasses hybridPC. In these cases, it approaches the performance level of BP-based models. These results are also illustrated in Figure 4C with example image reconstructions for each model. The superior representation learning capability of bPC over genPC suggests that a feedforward initialization through the bottom-up pathway in bPC serves as a fast amortised inference that nudges its initial neural activities towards the optimal state. This is particularly useful when inference steps are limited. The advantage of bPC over hybridPC is likely due to the active involvement of the bottom-up weights $V_{l}$ in the inference dynamics, which delivers sensory information to the latent neurons during iterative inference. This becomes more important when the sensory inputs are more complex, such as in the CIFAR datasets. The local nature of computations in bPC, which leads to lower learning efficiency but more plausible implementation, could underlie its slightly worse performance than BP-based models.
4.3 Combined supervised and unsupervised learning
Here, we integrate the supervised and unsupervised settings described previously to show that bPC can simultaneously develop discriminative capabilities and compressed representations. We trained bPC, hybridPC and their BP-based equivalents on the MNIST, Fashion-MNIST, and CIFAR-10 datasets. During training, we clamped layer $x_{1}$ to input images and partially clamped the top-layer $x_{L}$ to one-hot labels, while leaving the remaining neurons of $x_{L}$ free to learn complementary representations (see Figure 5A). Consistent with previous sections, MNIST and Fashion-MNIST models comprised two hidden layers of 256 neurons each, with layer $x_{L}$ composed of 40 neurons—10 dedicated to labels and 30 for representations. The CIFAR-10 model employed a convolutional architecture, with four hidden convolutional layers with max-pooling and a latent layer $x_{L}$ containing 266 neurons (10 for labels, 256 for representations). Additionally, activity decay was applied to the representational neurons in $x_{L}$. Classification accuracy was measured by providing only images (as in Section 4.1), while generative quality was assessed by reconstructing images from their representations (as in Section 4.2). The key difference is that images are reconstructed by clamping $x_{L}$ to the inferred representation and the image label.
Figure 5B reveals that bPC achieves classification accuracy on par with BP models across all datasets. Furthermore, it significantly surpasses hybridPC, particularly on CIFAR-10, where hybridPC exhibits more than 45% lower accuracy than bPC. In terms of generative quality, Figure 5D shows that bPC achieves reconstruction errors similar to hybridPC across all datasets. The reconstructed examples in Figure 5C illustrate that bPC goes beyond generating class-average images, capturing features such as shape, color, and spatial locations (e.g., it generates ‘4’s in different styles). However, fine details are absent in bPC’s CIFAR-10 reconstructions due to artefacts introduced by max-pooling operations in the discriminative bottom-up pathway, affecting the generative top-down reconstruction process. Consequently, bPC shows stronger grid-like artefacts than hybridPC, which uses only max-pooling for initialization, and substantially more than BP models with separate generative and discriminative layers. While removing max-pooling removes artefacts, it reduces classification accuracy (see SM). Nonetheless, even without max-pooling, bPC’s classification accuracy significantly outperforms hybridPC, while generative performance remains comparable. Future research could refine bPC’s architecture to balance classification accuracy with artefact-free generation.
Figure 5: bPC is the only PC model that can jointly learn low-dimensional representations of images and accurately classify them. A: Training set-up, where the latent layer is only partially clamped to class labels. B: Classification accuracy. C: Example reconstructions on MNIST and CIFAR10. D: Reconstruction RMSEs.
4.4 Shared latent layers prevent biased or overconfident models
To understand why bPC is able to simultaneously perform both discriminative and generative tasks well, we examine its energy landscape (Eq 4). We first consider the XOR task, a simple, $2$-dimensional nonlinear classification problem where the energy landscape can be visualised. We compare bPC’s energy landscape to those developed by individual genPC and discPC, to investigate the advantage of shared, rather than independent, latent layers. We trained bPC (Figure 6A bottom), genPC and discPC (Figure 6A top), each with two hidden layers of 16 neurons, by clamping $x_{1}$ to the $2$-dimensional inputs and $x_{4}$ to the scalar outputs. In this set-up, bPC has an equal number of parameters to that of genPC and discPC combined, but half the number of neurons. The resulting energy landscapes after training are shown in Figure 6B, where discPC exhibits broad regions of low energy, indicating overconfidence across a wide range of input configurations, including potentially OOD inputs. genPC, on the other hand, develops energy minima at the mean of each class, failing to capture the true structure of the XOR data. Notably, when discPC and genPC are combined into a single model with separate objectives, the resulting energy landscape retains both problems: broad and biased minima. In contrast, bPC learns sharp, well-localized minima around data points.
We then examined the generalisability of these observations, by evaluating the ability of the models considered in Section 4.1 to generate MNIST samples. Specifically, we clamped the topmost layer $x_{L}$ to a specific digit label, randomly initialized the input layer $x_{1}$, and performed inference updates until the model reached energy values equal or lower to the lowest 10% of equilibrium energies observed on MNIST test images. For a combined discPC-genPC model, we required the generated samples to satisfy this low-energy criterion for both components. This set-up serves as a sampling process, allowing us to inspect images that are considered as highly likely by the models. Figure 6C shows that samples generated by bPC are substantially more realistic than those from the combined discPC-genPC model, which often include excessive background noise and poorly defined digit shapes. To quantify sample quality, we used the Inception Score, that assesses how easily the generated digits can be labeled by a classifier [32], and the Fréchet Inception Distance (FID), which measures image quality and diversity [33]. bPC significantly outperformed the combined model, with higher Inception Scores (bPC: $6.05^{\pm 0.17}$ vs. combined: $3.62^{\pm 0.03}$, mean ± SEM over 5 seeds) and lower FID scores (bPC: $44.4^{\pm 2.2}$ vs. combined: $140.5^{\pm 2.1}$). We repeated the comparison with a BP-based combined model (discBP + genBP), and once again bPC produced higher-quality samples (see SM).
These results provide a detailed mechanistic explanation for the observations in Section 4.1. discPC performs poorly in image generation by confidently assigning low energy to nonsensical, OOD images with its broad energy minima. genPC recognizes only the mean image of each class as low-energy, thus relying on a template-matching mechanism where each new pattern is classified based on its distance to the mean class images. bPC, on the other hand, develops a more selective energy landscape as these two energy functions ($E_{disc}$ and $E_{gen}$) are jointly minimised and act as mutual regularisations on each other. This can be observed from bottom panel of Figure 6B, where $E_{gen}$ in bPC is ‘stretched’ by $E_{disc}$ to also include the individual data points, whereas $E_{disc}$ is ‘squeezed’ by $E_{gen}$ to have smaller, more dedicated minima.
Figure 6: bPC develops energy landscape suitable for both generation and classification. A: We compare bPC (bottom) to training a discPC and a GenPC separately (top). B: Visualisation of the models’ energy landscape after training on XOR. Shown are the discriminative, generative, and summed energies. C: MNIST samples considered highly likely (bottom 10% energy) by the models.
4.5 Learning in biologically relevant tasks
Figure 7: bPC performance in multimodal learning and robustness against missing information. A: Bimodal bPC and genPC. Note that bimodal bPC is fully equivalent to unimodal bPC in Section 4.1; it is bent here to serve visual purposes. B: Classification accuracy of bimodal PC models. C & D: Generation RMSE and examples. E: Model evaluation with partially occluded sensory inputs. F & G: Classification accuracy against percentage of missing pixels with MNIST and Fashion-MNIST respectively. H: Activity of $x_{1}$ after inference for genPC and bPC.
First, we examine bPC in multimodal learning, which is a typical way the brain forms associations [34]. When simulating supervised learning tasks, the information on the label is typically provided to a top layer for simplicity, however, when humans acquire information, the category is often represented in another modality. For example, when a child learns names of animals, these names are spoken by a parent and processed via the auditory cortex (Figure 7A). To assess multimodal capabilities, we compared bPC and genPC on a bimodal task using the MNIST and Fashion-MNIST datasets. We trained hierarchical models with one latent layer connected to two separate input layers (Figure 7A), one clamped to a one-hot vector of label and the other to the corresponding input image. The bPC model included bottom-up and top-down predictions between each input layer and the latent layer, whereas genPC incorporated only top-down predictions from the latent to the input layers.
After training, we evaluated cross-modal information transfer by measuring classification accuracy and average image reconstruction error. We provided input to one modality and ran iterative inference. Next, we compared the activity of the modality without input against either the correct label or the class-averaged image. Figure 7B, C and D illustrates that the bimodal bPC model significantly outperforms genPC in both classification accuracy and image reconstruction performance. This result is expected as bimodal bPC is equivalent to the bPC model introduced in Section 4.1 due to the bottom-up and top-down symmetry, which has been proven successful in both tasks. In other words, hidden layers of bPC naturally model associative cortical areas. On the other hand, genPC has to be purposely re-structured to model the multimodal association cortex, where one pathway predicts the image (similar to genPC) and the other predicts the label (similar to discPC). Consequently, bimodal genPC inherits limitations from both model types, resulting in poor classification and generation.
Next, we test the robustness of bPC to partially missing information, as the brain is known to retain important cognitive functions such as categorisation, even with incomplete or occluded visual scenes [35]. We used trained models (bPC, genPC, discPC, and discBP) from section 4.1 to assess their classification accuracy under increasing levels of missing pixels (from 0% to 90%) on MNIST and Fashion-MNIST. We fix the activity of observed neurons in $x_{1}$ to observed pixels and leave unobserved neurons free, initialized to zero activity as shown in figure 7E. To accommodate the slower inference convergence with missing data, we allow extended inference durations before evaluation.
As shown in Figure 7F and G, bPC maintains high classification accuracy even with up to 80% of pixels missing. genPC also retains reasonable performance, though its overall accuracy is lower than bPC across levels of occlusion. By contrast, discPC and discBP exhibit sharp performance drops beyond 50% missing input. bPC’s robustness stems from its top-down predictions: they actively fill in missing input through inference, reconstructing inputs by integrating top-down generative priors and bottom-up evidence. Figure 7H shows example input images where missing inputs are filled in by bPC. genPC shares this capacity but suffers from weak discriminative power. In contrast, discPC and discBP cannot infer missing information because they lack the top-down generative mechanism.
5 Conclusion
Inspired by empirical and theoretical insights into visual processing, we propose bidirectional predictive coding, a biologically plausible computational model of visual inference that integrates generative and discriminative processing. We demonstrate that bPC performs effectively across both supervised classification and unsupervised representation learning tasks, consistently outperforming or matching traditional predictive coding models. Our experiments reveal that the performance of bPC emerges from its ability to develop an energy landscape optimized simultaneously for both discriminative and generative tasks, thereby improving its robustness to out-of-distribution data. Furthermore, we show bPC’s effectiveness in biologically relevant scenarios such as multimodal integration and inference with partially missing inputs. Overall, bPC offers a hypothesis for how flexible inference could emerge in the brain, while also providing a method to enhance the robustness of discriminative models in machine learning applications.
Supplemental Material
The supplementary material is organised as follows:
We begin with a detailed explanation of the methods underlying the experiments presented in the main paper. Next, we provide summary tables containing the exact values plotted in the figures. We then explore the learning performance of three additional predictive coding models. Following that, we examine the effect of scaling the discriminative and generative energies in bPC on both supervised and unsupervised learning performance. We also analyse the impact of removing max-pooling on generative artefacts. Finally, we present samples generated by a combined discBP and genBP model.
Appendix A Details of experiments
Here, we detail the training setup and the evaluation procedure of the experiments in the paper. The code with all models and experiments will be available upon publication at https://github.com/Bogacz-Group. Our implementation of the predictive coding models was adapted from pcx_.
A.1 Datasets
We evaluate the models on four standard image classification benchmarks: MNIST [lecun2010mnist_], Fashion-MNIST [xiao2017fashionmnist_], CIFAR-10, and CIFAR-100 [krizhevsky2009learning_]. Below, we summarise their properties and how they were subdivided in our experiments for training, validation and testing. The images of all datasets were normalised so that pixel values varied between $-1$ and $1$.
  • MNIST: A dataset of grayscale handwritten digits (0–9).
  • Fashion-MNIST: A dataset of grayscale images of clothing items.
  • CIFAR-10: A dataset of colour images categorised into 10 classes of everyday objects.
  • CIFAR-100: Similar to CIFAR-10, but with 100 fine-grained object categories.
Table 1: Summary of dataset characteristics. The evaluation set of each dataset is split 50/50 into validation and test subsets.
Dataset Image Size Channels Classes Train Validation Test
MNIST $28\times 28$ 1 10 60,000 5,000 5,000
Fashion-MNIST $28\times 28$ 1 10 60,000 5,000 5,000
CIFAR-10 $32\times 32$ 3 10 50,000 5,000 5,000
CIFAR-100 $32\times 32$ 3 100 50,000 5,000 5,000
In addition to these benchmarks, we also consider the XOR learning task. XOR is a simple, non-linearly separable binary classification problem, often used to test the capacity of neural networks to learn non-linear functions. The input-output relationship of XOR is given in Table 2.
Table 2: XOR truth table with scaled inputs.
Input 1 Input 2 Output (XOR)
-1 -1 -1
-1 1 1
1 -1 1
1 1 -1
A.2 Model architectures
We consider five different neural architectures in our experiments, selected based on the dataset and the learning setting.
MLP for MNIST and Fashion-MNIST For MNIST and Fashion-MNIST in Sections 4.1, 4.2 and 4.3, we use a network with neural layers of dimension [784, 256, 256, latent_size]. The latent layer size is set to 10 for supervised experiments, 30 for unsupervised experiments, and 40 for experiments with partial clamping of the top-layer activity ($x_{4}$). Both generative and discriminative predictions between layers follow the form $f(Wx+b)$, where $f$ is a nonlinearity, $W$ is a weight matrix, $b$ is a bias vector, and $x$ is the activation of the previous layer. This is the transformation of a layer in a multilayered perceptron (MLP). Consequently, unidirectional models (discPC, genPC, discBP, genBP) have the architecture of an MLP. hybridPC and bPC have the same number of parameters as two MLPs (one generative and one discriminative).
CNN for unsupervised learning with CIFAR-10 and CIFAR-100 In the unsupervised representational learning experiments (Section 4.2) on CIFAR-10 and CIFAR-100, we use a model architecture with convolutions and a representational layer $x_{L}=x_{6}$ of 256 neurons. In this model, discriminative predictions consist of 5 strided convolution layers, with a stride of two, and one MLP layer. Generative predictions mirror the discriminative layers with 5 transposed convolution layers, with a stride of two, and one MLP layer. The convolutional-based layers perform the transformation $f(c(x))$, where $c$ is a (transposed) convolution layer and $f$ is a nonlinearity. A full description of the convolution-based layers can be found in Table 3.
VGG-5 for combined supervised and unsupervised learning with CIFAR-10 For the CIFAR-10 experiments that combine supervised and unsupervised representational learning in Section 4.3, we use a VGG-5-style network with a 266-dimensional latent layer $x_{L}=x_{6}$ (10 neurons for class labels and 256 for representations). In this model, discriminative predictions consist of 4 convolutional layers with max pooling and one MLP layer. The generative predictions consist of 4 transposed convolution layers and one MLP layer. The discriminative convolution layers have a stride of one, and the transpose convolutions have a stride of two to compensate for the dimensionality reduction of the max-pooling. A full description of the convolution-based layers can be found in Table 3.
For the above three models, the activation function used in the discriminative prediction from $x_{L-1}$ to $x_{L}$ is an identity layer (no activation). The generative prediction from layer $x_{2}$ to $x_{1}$ uses a $tanh(\cdot)$ activation to constrain outputs to the range [-1, 1], matching normalised pixel values. The other activation functions are the same for the whole network and are determined using a hyperparameter search.
MLP for XOR Task For the XOR task in Section 4.4, we use a network with neural layers of dimension $[2,16,16,1]$. The generative and discriminative predictions are equivalent to the ones used for MNIST and Fashion-MNIST, with the exception that: a sigmoid activation is used for the prediction from $x_{3}$ to $x_{4}$ and an identity activation is used for the prediction from $x_{2}$ to $x_{1}$. These changes stabilise training and help avoid bias because the tanh activation used for (Fashion-)MNIST saturates near ±1, but never reaches these values, causing the model to systematically under-predict XOR inputs.
MLP for Bimodal Learning Task In the bimodal generation task of Section 4.5, the network consists of two input layers, with 784 (image) and 10 (label) neurons, and one latent layer with 256 neurons. The predictions between the latent layer and the input layers are MLP layers. The bimodal bPC model is equivalent to the bPC model trained on (Fashion-)MNIST in Section 4.1, except it only has one hidden layer. In contrast, the bimodal genPC is different from the models considered in Section 4.1, and it can be re-interpreted as a unidirectional model with one generative MLP layer and one discriminative MLP layer. For fair comparison, we include an additional nonlinearity in the prediction layers of the genPC variant. The MLP layers become $f_{1}(Wf_{2}(x)+b)$. For the image modality, $f_{1}=\tanh(\cdot)$; for the label modality, $f_{1}$ is the identity. The additional nonlinearity, $f_{2}$, ensures parity with bPC models, which include nonlinear discriminative projections.
Table 3: Detailed architectures of convolution-based models. Convolution and transposed convolution have the same kernel sizes and paddings. The values in the brackets give the parameter value for each convolution-based layer in the model, starting from the layer closest to the input image.
CNN VGG-5
Channel Sizes [32, 64, 128, 256, 512] [128, 256, 512, 512]
Kernel Sizes [3, 3, 3, 3, 3] [3, 3, 3, 3]
Paddings [1, 1, 1, 1,1] [1, 1, 1, 1]
Strides conv. [2, 2, 2, 2,2] [1, 1, 1, 1]
Strides transposed conv. [2, 2, 2, 2,2] [2, 2, 2, 2]
Output Paddings transposed conv. [1, 1, 1, 1,1] [0, 0, 0, 0]
Pooling Window - $2\times 2$
Pooling Stride - 2
A.3 Training procedures
Require : Model with neural activities $x$, parameters $\theta$, energy $E$ and initialisation function $init(\cdot)$. Dataset $\{y_{p}\}_{p=1}^{P}$ with $P$ mini-batches of $B$ elements. Number of epochs $N$. Activity optimiser $optim_{x}(\cdot)$, number of activity updates $K$, and parameter optimiser $optim_{\theta}(\cdot)$
for $n=1$ to $N$ do
       for $p=1$ to $P$ do
             // Independent inference for each sample in batch
             $x_{b}\leftarrow init(y_{p,b}),\quad 1\leq b\leq B$
             for $k=1$ to $K$ do
                   $x_{b}\leftarrow optim_{x}(\frac{\partial E_{b}}{\partial x_{b}}),\quad 1\leq b\leq B$
            // Sum of parameter updates for batch
             $\theta\leftarrow optim_{\theta}(\frac{1}{B}\sum_{b}^{B}\frac{\partial E_{b}}{\partial\theta})$
Algorithm 1 Training procedure of PC models.
Require : Model with parameters $\theta$ and forward pass $fp(\cdot,\theta)$. Energy function $E$. Dataset $\{y_{p}\}_{p=1}^{P}$ with $P$ mini-batches of $B$ elements. Number of epochs $N$. Parameter optimiser $optim_{\theta}(\cdot)$
for $n=1$ to $N$ do
       for $p=1$ to $P$ do
             // Independent forward pass for each sample in batch
             $\hat{y}_{b}\leftarrow fp(y_{p,b},\theta),\quad 1\leq b\leq B$
             // Sum of parameter updates for batch
             $\theta\leftarrow optim_{\theta}(\frac{1}{B}\sum_{b}^{B}\frac{\partial E(\hat{y}_{b},y_{p,b})}{\partial\theta})$
Algorithm 2 Training procedure of BP models.
We followed the procedures outlined in Algorithm 1 to train predictive coding models and hybridBP, and the procedure outlined in Algorithm 2 for the remaining backpropagation models. For a given task, both algorithms were trained on the same dataset for the same number of epochs.
Optimisers
The optimisation procedures differ between activity updates and parameter updates. In Algorithm 1, neural activities are updated using stochastic gradient descent with momentum. This optimiser outperforms Adam in PC models [pcx_]. In contrast, Algorithm 2 does not include activity updates and only optimises parameters. For parameter updates in both algorithms, we use the AdamW optimiser [loshchilov2017decoupled_]. The optimiser hyperparameters were determined via a hyperparameter search.
Initialisation functions
The initialisation procedure in Algorithm 1 generally consists of (1) clamping the activity $x_{1}$ and $x_{L}$ to images and labels depending on the learning task and (2) performing a feedforward sweep. This sweep initialises the activity of each layer by propagating inputs through the model in the direction of prediction. For example, a bottom-up feedforward sweep for a three-layer model initialises the second and third layers as: $x_{2}=f(x_{1})$ and $x_{3}=V_{2}f(V_{1}f(x_{1}))$. This approach, originally proposed by whittington2017approximation_, significantly reduces the time required for inference and improves learning. The following models use bottom-up feedforward initialisation: discPC, hybridPC, bPC, and hybridBP. The genPC model, by contrast, uses a top-down feedforward initialisation. The difference between hybridBP and the other models initialised using a bottom-up sweep is that hybridBP contains a single layer of neural activities, $x_{L}$, that can be initialised using a full bottom-up forward pass.
The bimodal genPC model from Section 4.5 differs from the other PC models in its initialisation procedure. Because there are no direct prediction paths from the two modality-specific input layers to the shared latent layer, a standard feedforward sweep cannot be used to initialise the latent layer. Instead, we tested two alternatives for initialising the latent layer: zero-initialisation and Xavier uniform initialisation. The best initialisation depended on the task and was determined with a hyperparameter search. For the bimodal bPC model, we retained the feedforward sweep used for other bPC models, using a forward sweep from the image modality to the latent layer.
Energy functions
The energy functions for all PC and BP models are listed in Table 4. For PC models, the energy functions follow the formulations described in the main text, except that we use a generalised notation: $f_{l}$ denotes the transformation applied at layer $l$, allowing for models beyond standard MLPs. BP model energy functions use squared errors to remain consistent with PC formulations.
The hybridBP model includes a stop-gradient operation in its energy function. This ensures that its second loss term is only applied during parameter updates, not iterative inference. This loss term objective is to improve the discriminative initialisation of the model’s latent layer. In the broader machine learning literature, hybridBP can be viewed as a generative model that performs Bayesian inference by combining amortised inference with iterative refinement of the latent variables [marino2018iterative_].
In the case of the autoencoder (AE), the energy function departs from standard autoencoders to accommodate experiments described in Section 4.3, where part of the latent state $x_{L}$ is fixed to class labels. The AE energy includes an additional term: $\|x_{L,ae}-f_{p,disc}(x_{1},V)\|_{2}^{2}$ where $x_{L,ae}:=[x_{L,0:k-1},f_{p,disc}(x_{1},V)_{k:D}]$ that combines $k$ fixed neurons and discriminative predictions. This term penalises the error between the fixed components of $x_{L,ae}$ and their prediction from the discriminative part of the model. Notably, this term is only non-zero for the fixed neurons. The additional energy term is also equal to zero in the representational learning experiments of Section 4.2 because none of the neurons of $x_{L,ae}$ are fixed. For reference, the discriminative part of our AE is usually referred to as the encoder and the generative part as the decoder.
In the combined supervised and unsupervised learning task (Section 4.3), we apply separate energy scaling to different parts of the models. Unsupervised learning benefits from equal weighting of discriminative and generative energies, while supervised tasks perform better when the generative energy is down-scaled. To balance these, we set $\alpha_{disc}$ of the free (unclamped) neurons in $x_{L}$ to $\alpha_{gen}$. This allows for the generative energy to be down-scaled in the whole network except for the free neurons of $x_{L}$ that specialise in unsupervised representational learning. This adjustment was used for bPC, hybridBP, and AE, but not for hybridPC. In hybridPC, the discriminative weights predicting free and fixed neurons are independent due to the use of local learning rules. Moreover, these discriminative weights have minimal influence on the generative weights, as inference is fully driven by the generative energy. As a result, scaling the discriminative and generative energies in different parts of the model merely scales the corresponding gradients. This effect has no impact on learning because the AdamW optimiser normalises gradients. In contrast, in hybridBP, some parameters are shared in predicting both free and fixed neurons. This introduces a dependency between the scaling of energies associated with free and fixed neurons in $x_{L}$. Therefore, we introduced explicit scaling factors $\alpha_{disc}$ and $\alpha_{gen}$ in hybridBP’s energy function for this task.
Model Loss Function
discPC $E_{\text{disc}}(x,V)=\sum\nolimits_{l=2}^{L}\frac{1}{2}\|x_{l}-f_{l}(x_{l-1},V_{l-1})\|_{2}^{2}$
discBP $\mathcal{L}_{disc}(x_{1},x_{L},V)=\frac{1}{2}\|x_{L}-f_{p,disc}(x_{1},V)\|_{2}^{2}$
genPC $E_{\text{gen}}(x,W)=\sum\nolimits_{l=1}^{L-1}\frac{1}{2}\|x_{l}-f_{l}(x_{l+1},W_{l+1})\|_{2}^{2}$
genBP $\mathcal{L}_{gen}(x_{1},x_{L},W)=\frac{1}{2}\|x_{1}-f_{p,gen}(x_{L},W)\|_{2}^{2}$
hybridPC $E_{\text{hybrid}}(x,W,V)=\sum\nolimits_{l=1}^{L-1}\frac{1}{2}\|x_{l}-f_{l}(x_{l+1},W_{l+1})\|_{2}^{2}+\sum\nolimits_{l=2}^{L}\frac{1}{2}\|\textrm{sg}(x_{l})-f_{l}(\textrm{sg}(x_{l-1}),V_{l-1})\|_{2}^{2}$
hybridBP $\mathcal{L}_{\text{hybrid}}(x_{1},x_{L},W,V)=\frac{1}{2}\|x_{1}-f_{p,gen}(x_{L},W)\|_{2}^{2}+\frac{1}{2}\|\textrm{sg}(x_{L})-f_{p,disc}(x_{1},V)\|_{2}^{2}$
bPC $E(x,W,V)=\sum\nolimits_{l=1}^{L-1}\frac{\alpha_{\text{gen}}}{2}\|x_{l}-f(x_{l+1},W_{l+1})\|_{2}^{2}+\sum\nolimits_{l=2}^{L}\frac{\alpha_{\text{disc}}}{2}\|x_{l}-f(x_{l-1},V_{l-1})\|_{2}^{2}$
AE $\mathcal{L}(x_{1},x_{L},W,V)=\frac{\alpha_{\text{gen}}}{2}\|x_{1}-f_{p,gen}(x_{L,ae},W)\|_{2}^{2}+\frac{\alpha_{\text{disc}}}{2}\|x_{L,ae}-f_{p,disc}(x_{1},V)\|_{2}^{2},$
        $x_{L,ae}:=[x_{L,0:k-1},f_{p,disc}(x_{1},V)_{k:D}]$
Table 4: Energy functions for the predictive coding and backpropagation models considered in the paper. In the PC models, $f_{l}(x,\theta)$ is the transformation between layers with input $x$ and parameters $\theta$. In the BP models, $f_{p}$ denotes a forward pass and $x_{L,ae}$ is an activity vector composed of $k$ fixed neurons and fills the remaining outputs of the discriminative forward pass of the AE.
A.4 Training hyperparameters
In this Section, we report the hyperparameter search space for the experiments in Sections 4.1, 4.2, 4.3, and 4.5, as well as the training parameters for the models in Section 4.4. All hyperparameter tuning was performed using Bayesian optimisation from Weights and Biases [wandb_].
Tables 5, 6, and 7 list the hyperparameter search spaces for the models trained in Sections 4.1, 4.2, and 4.3, respectively. The hyperparameter search was conducted using the validation sets of each dataset. In Section 4.1, discPC and genPC were tuned for classification accuracy and generation RMSE separately, while bPC and hybridPC were jointly tuned on both metrics. This ensures a fair comparison, as bPC and hybridPC have twice as many parameters as discPC and genPC. For joint tuning, we combined the two metrics using the objective $2\cdot(1-\text{accuracy}/100)+\text{RMSE}$. In Section 4.2, the models were tuned for the image reconstruction MSE, after which the RMSE was reported for consistency with the generation RMSE. In Section 4.3, all models were jointly optimised for classification accuracy and reconstruction MSE. We combined the two metrics using $2\cdot(1-\text{accuracy}/100)+\text{MSE}$ for MNIST and Fashion-MNIST, and $(1-\text{accuracy}/100)+4\cdot\text{MSE}$ for CIFAR-10. The RMSE was also reported after tuning instead of the MSE. The weightings in the combined objectives compensate for the scale difference between the metrics.
Table 8 reports the training parameters for the models used to generate MNIST samples in Section 4.4. The leaky ReLU activation function was used because it generated the best samples across model types, and the other parameters were set to default values that ensured stable learning across models.
Table 9 lists the hyperparameter search space for the bimodal bPC and genPC models trained in Section 4.5. The genPC model was tuned separately for classification accuracy and reconstruction RMSE, while the bPC model was tuned jointly on both metrics. The joint tuning followed the same combined metric as described above for MNIST. During training and evaluation, more inference steps were used for genPC than bPC, as genPC lacks an effective feedforward initialisation scheme, which slows inference convergence.
Table 5: Hyperparameter search configuration for experiments in Section 4.1 for both MNIST and Fashion-MNIST datasets.
Parameter bPC hybridPC discPC genPC genBP discBP
Epoch 25
Batch Size 256
Activation [leaky relu, tanh, gelu]
$lr_{x}$ (1e-3, 5e-1)2 (1e-4, 5e-2)2 -
$\text{momentum}_{x}$ [0.0, 0.5, 0.9] -
$lr_{\theta}$ (1e-5, 1e-3)2 (1e-6, 1e-4)2 (1e-5, 1e-3)2
[…]
T 8 -
T eval 100 -
$\alpha_{gen}$ 1e[0,-4;-1]1 - -
$\alpha_{disc}$ 1 - -
1: “[a, b; c]” denotes a sequence of values from a to b with a step size of c.
2: “(a, b)” represents a log-uniform distribution between a and b.
Table 6: Hyperparameter search configuration for experiments in Section 4.2 for MNIST , Fashion-MNIST, CIFAR-10, and CIFAR-100 datasets. Parameters unique to the models trained on MNIST and Fashion-MNIST are indicated with MLP. Parameters unique to the CIFAR datasets are indicated with CNN.
Parameter bPC genPC hybridPC hybridBP AE
Epoch (MLP) 25
Epoch (CNN) 50
Batch Size 256
Activation [leaky relu, tanh, gelu]
$lr_{x}$ (1e-3, 5e-1)1 -
$\text{momentum}_{x}$ [0.0, 0.5, 0.9] -
$lr_{\theta}$ (MLP) (1e-5, 1e-3)1
$lr_{\theta}$ (CNN) (1e-5, 1e-3)1,2
[…]
T (MLP) 8 -
T (CNN) 32 -
T eval 100 -
$\alpha_{gen}$ 1 - - 1
$\alpha_{disc}$ 1 - - 1
1: “(a, b)” represents a log-uniform distribution between a and b.
2 Learning rates of $\theta$ were scaled with warmup-cosine-annealing scheduler without restart.
Table 7: Hyperparameter search configuration for experiments in Section 4.3. Parameters unique to the models trained on MNIST and Fashion-MNIST are indicated with MLP. Parameters unique to the CIFAR-10 dataset are indicated with VGG. Two activity optimisers and two parameter optimisers are used for the VGG models. This allows the models to have different learning rates for different task-specific parts of the networks. One activity optimiser was used for the free neurons of $x_{L}$ with learning rate $lr_{x,\mathrm{free}}$, and another for all other neurons with learning rate $lr_{x}$. Both activity optimisers use the same momentum parameter. One parameter optimiser with learning rate $lr_{\theta,gen}$ was used for the generative parameters and the discriminative parameters predicting the free neurons of $x_{L}$. The other parameter optimiser was used for the remaining discriminative parameters. Both parameter optimisers use the same weight decay.
Parameter bPC hybridPC hybridBP AE
Epoch (MLP) 25
Epoch (VGG) 50
Batch Size 256
Activation [leaky relu, tanh, gelu]
$lr_{x}$ (1e-3, 5e-1)1 -
$lr_{x,free}$ (VGG) (1e-3, 5e-1)1 -
$\text{momentum}_{x}$ [0.0, 0.5, 0.9] -
$lr_{\theta}$ (MLP) (1e-5, 1e-3)1
$lr_{\theta,disc}$ (VGG) (1e-5, 1e-3)1,3
$lr_{\theta,gen}$ (VGG) (1e-5, 1e-2)1,3
[…]
T (MLP) 8 -
T (VGG) 32 -
T eval 100 -
$\alpha_{gen}$ (MLP) 1e[0,-4;-1]1 - 1e[0,-4;-1]1 1e[0,-4;-1]1
$\alpha_{gen}$ (VGG) 1e[-4,-7;-1]1 - 1e[0,-8;-1]1 1e[0,-8;-1]1
$\alpha_{disc}$ 1 - 1 1
1: “(a, b)” represents a log-uniform distribution between a and b.
2: “[a, b; c]” denotes a sequence of values from a to b with a step size of c.
3 Learning rates of $\theta$ were scaled with warmup-cosine-annealing scheduler without restart.
Table 8: Hyperparameters for models used to sample MNIST images in Section 4.1.
Parameter bPC genPC discPC
Epoch 25
Batch Size 256
Activation leaky relu
$lr_{x}$ 0.01
$\text{momentum}_{x}$ 0.0
$lr_{\theta}$ 1e-4
[…]
T 8
T eval 100
$\alpha_{gen}$ 1e-4 -
$\alpha_{disc}$ 1 -
Table 9: Hyperparameter search configuration for bimodal models in Section 4.5 for both MNIST and Fashion-MNIST datasets.
Parameter bPC genPC
Epoch 25
Batch Size 256
Activation [leaky relu, tanh, gelu]
$lr_{x}$ (1e-3, 5e-1)2
$momentum_{x}$ [0.0, 0.5, 0.9]
$lr_{\theta}$ (1e-5, 1e-3)2
$weight\_decay_{\theta}$ (1e-5, 1e-2)2
T 8 20
T eval 100 1000
$\alpha_{gen}$ 1e[0,-4;-1]1 -
$\alpha_{disc}$ 1 -
1: “[a, b; c]” denotes a sequence of values from a to b with a step size of c.
2: “(a, b)” represents a log-uniform distribution between a and b.
A.5 Evaluation
In our experiments, we consider six different evaluation procedures discussed below.
Classification accuracy
We evaluate classification performance using accuracy, defined as the percentage of input images correctly classified by each model.
For the PC models, classification is performed by first setting $x_{1}$ to the input image. We initialise the remaining layers using a bottom-up feedforward sweep, followed by 100 steps of iterative inference. The predicted class is determined by identifying the neuron in the output layer $x_{L}$ (corresponding to class encoding units) with the highest activity. The bottom-up initialisation significantly reduces the time required to reach steady state during inference. However, bottom-up initialisation does not apply to genPC and bimodal genPC due to the absence of bottom-up predictive pathways. In genPC, we initialise $x_{L}$ to zero and perform a top-down feedforward sweep instead. For bimodal genPC, the latent layer is initialised identically to the procedure used during training.
For the BP models, classification is performed via a standard discriminative forward pass from the input image, with the predicted class taken as the one corresponding to the highest output activation.
Conditional generation of mean class images
We evaluate the generative performance of the models by computing the root mean squared error (RMSE) between images generated for each class and the corresponding class-average image. We obtain the class-average image by computing the average image across all images belonging to that class in the evaluation set. To compute the RMSE, we calculate the squared error for each pixel between the generated image and the average image for the corresponding class. These errors are averaged across all pixels and classes, and the square root of this mean is reported as the final RMSE.
To generate an image from a PC model conditioned on a class label, we set the top-layer activity $x_{L}$ to the one-hot encoding of the target label. The remaining layers are then initialised using a top-down feedforward sweep. One exception is the discPC model, which does not support top-down initialisation. For discPC, we initialise the input layer $x_{1}$ to zero activity and then perform a bottom-up feedforward sweep. Following initialisation, we run 100 steps of iterative inference and extract the final activity of the input layer $x_{1}$ as the generated image.
For the BP models, we generate an image by performing a generative forward pass starting from a one-hot label vector.
Image reconstruction from representations
We assess the quality of learned representations by measuring the RMSE between reconstructed images and their original inputs. RMSE is computed in the same manner as for conditional image generation.
For PC models and hybridBP, the reconstruction procedure is as follows: (1) clamp $x_{1}$ to the input image, (2) initialise the remaining layers via a bottom-up feedforward sweep, (3) perform 100 steps of iterative inference, (4) clamp $x_{L}$ to its activity after the inference, (5) re-initialise the other layers using a top-down feedforward sweep, (6) run 100 additional inference steps, and (8) record the final activity in $x_{1}$ as the reconstructed image.
For the autoencoder, an image is reconstructed by performing a discriminative forward pass using the input image to obtain a representation, followed by a generative forward pass from the representation to get the reconstructed image.
In Section 4.3, where labels are also provided during reconstruction, the above procedure is slightly modified. For PC models and hybridBP, certain neurons of $x_{L}$ are additionally clamped to the ground truth label during the first step. During the fourth step, the label components and the remaining latent representation (recorded after inference) are clamped in $x_{L}$. For the AE, we retain the standard discriminative forward pass, but replace the predicted label portion of its representation with the true label before passing it through the generative model to produce the reconstructed image.
Visualisation of Energy Landscapes on XOR
We visualise the energy landscapes of trained genPC, discPC, and bPC models on the XOR task. To generate the landscape, we clamp the input layer neurons to a 2D coordinate within the range [-3,3], sampled at intervals of 0.25 along both axes. Simultaneously, we clamp the output layer $x_{L}$ to the one-hot encoding of one of the two classes. For each coordinate–label combination, we run 10,000 steps of iterative inference to ensure convergence to equilibrium and record the final energy of the model. This procedure is repeated over the grid of 2D inputs and both class labels, allowing us to plot the full energy landscape. For fair comparison, we also include the combined energy landscape of discPC and genPC. Together, the combined parameter count of discPC and genPC matches the total parameter count of bPC. Equal weighting is applied to the energy values of the discPC and genPC models when adding their energy, as we use equal weighting for the discriminative and generative components in the bPC model ($\alpha_{disc}=\alpha_{gen}=1$).
MNIST image generation with PC models
We evaluate the generative capabilities of the bPC model and a combined genPC and discPC model on the MNIST dataset. The goal is to assess how likely these models are to assign low energy to implausible, label-inconsistent samples.
As a baseline, we first estimate the energy distribution over the test set for each model. For each test image, we clamp $x_{1}$ to the image and $x_{L}$ to its associated label. The model is then initialised as during training, and we run 50,000 steps of iterative inference to reach a steady state. We compute the 10th percentile from the resulting energy values as an energy threshold for high-quality, in-distribution samples. This process was repeated independently for bPC, genPC and discPC.
To generate images conditioned on a label, we use the following procedure: (1) randomly initialise the input layer $x_{1}$ by sampling each neuron’s activity uniformly from the interval [-1,1], (2) clamp $x_{L}$ to the target label, (3) initialise all hidden layers to zero, (4) run iterative inference until the median energy of the batch falls below the previously determined 10th percentile threshold, (5) retain the 50% of samples within the batch that fall below or equal to the energy threshold and discard the rest. Using this method, we generate 256 samples per model.
In the combined genPC and discPC model, the input layer $x_{1}$ is shared by both models, and its updates during inference are influenced by both energy functions. However, genPC typically exhibits energy magnitudes significantly larger than those of discPC. To prevent one component from dominating the updates to $x_{1}$, we scale the generative energy by a factor equal to the ratio of the discPC energy threshold to the genPC energy threshold. Additionally, inference is only terminated once the generative and the discriminative energies fall below their respective thresholds.
The models used for sample generation are trained separately from those used in the supervised experiments from Section 4.1. During preliminary testing, we found that generative quality strongly depends on the choice of activation function. In particular, leaky-ReLU yielded superior image samples. As a result, we train all models using leaky-ReLU activations and a shared set of training hyperparameters. These parameters are selected for their stability and effectiveness across all model types for classification and generation tasks.
We assess the quality and diversity of the generated samples using two standard metrics:
  • Fréchet Inception Distance (FID): The FID measures the distance between the distributions of real and generated images [HeuselRUNKH17_]. We use the public implementation of FID from Seitzer2020FID_, but modify it to use a ResNet classifier trained on MNIST, instead of the original Inception model trained on ImageNet. This ensures that the FID score better reflects visual quality and diversity on the MNIST domain.
  • Inception Score (IS): The Inception Score evaluates how easily a classifier can identify the class of a generated image [salimans2016improvedtechniquestraininggans_]. We used a publicly available implementation for the MNIST dataset [inception_].
Classification accuracy with partially missing inputs
To assess model robustness to missing input data, we evaluate classification accuracy under varying levels of input occlusion. This evaluation follows the same procedure as standard classification (described above), but for images missing a random subset of pixels. The proportion of missing pixels ranges from 10% to 90%, in increments of 10%. Missing pixels are selected uniformly at random, independently of their spatial location within the image. For all models, missing pixels in $x_{1}$ are initialised to zero activity and are left unclamped, allowing the model to update their values during inference. We repeat the same classification procedure as before, but we increased the number of inference steps to 600,000 to ensure convergence. This is necessary because predictive coding models converge more slowly when input information is incomplete [tang2023recurrent_]. In addition to measuring classification accuracy, we also record the post-inference activity of neurons in $x_{1}$ to qualitatively assess the model’s ability to fill in the missing input.
A.6 Compute resources
All experiments were conducted on NVIDIA RTX A6000 GPUs. Training an MLP model for MNIST and Fashion-MNIST experiments across tasks took less than one minute. Training unsupervised learning models on CIFAR-10 and CIFAR-100 of Section 4.2 took approximately 15 minutes. Training the combined supervised and unsupervised models of Section 4.3 took approximately one hour and 15 minutes. The majority of the compute was spent on hyperparameter tuning. The total training time for hyperparameter tuning of the models of Section 4.1 is ±$\pm 50$h. The total training time for hyperparameter tuning of the models of Section 4.2 is ±$\pm 30$h for MNIST and Fashion-MNIST and ±$\pm 400$h for CIFAR-10 and CIFAR-100. The total training time for hyperparameter tuning of the models of Section 4.3 is ±$\pm 25$h for MNIST and Fashion-MNIST and ±$\pm 750$h for CIFAR-10. The total training time for the tuning of the bimodal models of Section 4.5 is ±$\pm 20$h.
Appendix B Results
Tables 10 to 12 report the result illustrated in Figures 3, 4, 5, and 7. These results are obtained on the test set of the datasets for five different weight initialisations.
Table 10: Classification accuracy and class average image generation for models considered in Section 4.1. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
Acc % RMSE
Model MNIST Fashion-MNIST MNIST Fashion-MNIST
bPC $98.10^{\pm 0.05}$ $89.24^{\pm 0.12}$ $0.0581^{\pm 0.0004}$ $0.0415^{\pm 0.0005}$
hybridPC $86.22^{\pm 0.15}$ $80.34^{\pm 0.11}$ $0.0612^{\pm 0.0003}$ $0.0480^{\pm 0.0016}$
genPC $83.48^{\pm 0.21}$ $77.00^{\pm 0.16}$ $\mathbf{0.0198^{\pm 0.0001}}$ $0.0140^{\pm 0.0001}$
discPC $98.43^{\pm 0.01}$ $\mathbf{89.74^{\pm 0.14}}$ $0.3133^{\pm 0.0224}$ $0.3326^{\pm 0.0024}$
BP $\mathbf{98.48^{\pm 0.10}}$ $89.66^{\pm 0.11}$ $\mathbf{0.0198^{\pm 0.0001}}$ $\mathbf{0.0128^{\pm 0.0001}}$
Table 11: Image reconstruction RMSE from latent representations for models considered in Section 4.2. Lower RMSE is better. We report the mean +/- sem over five seeds.
Model MNIST Fashion-MNIST CIFAR-10 CIFAR-100
bPC $0.2320^{\pm 0.0010}$ $0.2497^{\pm 0.0004}$ $0.1311^{\pm 0.0005}$ $0.1366^{\pm 0.0007}$
genPC $0.2473^{\pm 0.0020}$ $0.2868^{\pm 0.0013}$ $0.1837^{\pm 0.0009}$ $0.2077^{\pm 0.0003}$
hybridPC $0.2401^{\pm 0.0012}$ $0.2508^{\pm 0.0007}$ $0.1664^{\pm 0.0015}$ $0.2089^{\pm 0.0071}$
AE $\mathbf{0.1565^{\pm 0.0006}}$ $\mathbf{0.1868^{\pm 0.0001}}$ $0.1135^{\pm 0.0050}$ $0.1171^{\pm 0.0042}$
BP $0.1969^{\pm 0.0004}$ $0.2084^{\pm 0.0002}$ $\mathbf{0.0964^{\pm 0.0004}}$ $\mathbf{0.0983^{\pm 0.0002}}$
Table 12: Classification accuracy and image reconstruction RMSE from latent representations for models considered in Section 4.3. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
Acc % RMSE
Model MNIST Fashion-MNIST CIFAR-10 MNIST Fashion-MNIST CIFAR-10
bPC $97.06^{\pm 0.08}$ $88.24^{\pm 0.08}$ $85.06^{\pm 0.11}$ $0.2394^{\pm 0.0002}$ $0.2531^{\pm 0.0022}$ $0.3036^{\pm 0.0014}$
hybridPC $86.15^{\pm 0.36}$ $80.11^{\pm 0.11}$ $37.20^{\pm 0.43}$ $0.2418^{\pm 0.0006}$ $0.2527^{\pm 0.0003}$ $0.3003^{\pm 0.0051}$
AE $98.32^{\pm 0.06}$ $89.49^{\pm 0.10}$ $89.97^{\pm 0.12}$ $0.1576^{\pm 0.0005}$ $0.2164^{\pm 0.0003}$ $0.1599^{\pm 0.0002}$
hybridBP $\mathbf{98.48^{\pm 0.03}}$ $\mathbf{89.61^{\pm 0.13}}$ $\mathbf{90.11^{\pm 0.11}}$ $\mathbf{0.1249^{\pm 0.0004}}$ $\mathbf{0.1707^{\pm 0.0004}}$ $\mathbf{0.1048^{\pm 0.0011}}$
Table 13: Classification accuracy and class average image generation for bimodal genPC and bPC considered in Section 4.5. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
Acc % RMSE
Model MNIST Fashion-MNIST MNIST Fashion-MNIST
bPC $\mathbf{97.80^{\pm 0.05}}$ $\mathbf{89.05^{\pm 0.10}}$ $\mathbf{0.0506^{\pm 0.0007}}$ $\mathbf{0.0431^{\pm 0.0004}}$
genPC $88.28^{\pm 0.10}$ $80.53^{\pm 0.14}$ $0.1561^{\pm 0.0004}$ $0.2027^{\pm 0.0031}$
Appendix C Other types of predictive coding models
In this experiment, we train three additional types of predictive coding models on the supervised task described in Section 4.1: (1) PC along arbitrary graphs [arbitrary_graph_], (2) bPC with shared weights for discriminative and generative predictions [qiu2023deep_], and (3) discPC with activity decay during generation [Sun2020APN_].
PC along arbitrary graphs (agPC) differs from bPC in that each neuron has a single energy function, with predictions computed from all incoming connections. In contrast, bPC uses separate energy terms for bottom-up (discriminative) and top-down (generative) predictions. In this formulation, the energy associated with an agPC layer $x_{l}$ is given by:
$E_{l}=\frac{1}{2}\left\|x_{l}-f(Wx_{l+1}+Vx_{l-1}+b)\right\|_{2}^{2},$
where $W$ and $V$ are top-down and bottom-up weights, respectively. We use the same training algorithm, hyperparameter tuning, and evaluation procedure as bPC, modifying only the energy function. Initialisation is done using bottom-up sweeps (for training and classification evaluation) and top-down forward sweeps (for generation evaluation), matching bPC. We also tested zero and Xavier initialisation. However, this resulted in worse learning performance.
bPC with shared weights (shared bPC) uses a single set of weights for both discriminative and generative predictions. Its energy function associated with a layer $x_{l}$ is given by:
$E_{l}=\frac{\alpha_{\text{gen}}}{2}\left\|x_{l}-f(W_{l+1}x_{l+1})\right\|_{2}^{2}+\frac{\alpha_{\text{disc}}}{2}\left\|x_{l}-f(W_{l}^{\top}x_{l-1})\right\|_{2}^{2},$
where $W_{l}^{\top}$ is reused for both directions. This model omits bias terms and relies on non-local computations. We follow the same training, tuning, and evaluation protocol as bPC.
discPC with activity decay (decay discPC) extends standard discPC by adding an activity decay term during generation. We use the same training and hyperparameter tuning protocol as discPC. Additionally, we tune an activity decay rate from a log-uniform distribution over the range $[10^{-5},1]$. During generative evaluation, we increase the number of inference steps to 10,000 due to slower convergence and we initialise neurons to zero activity before generation for consistency with Sun2020APN_.
Table 14 reports classification accuracy and reconstruction RMSE for these models. Figure 8 illustrates the images generated by each model for each MNIST and Fashion-MNIST class. None of the models match bPC in both classification and generative performance.
Table 14: Classification accuracy and class average image generation for agPC, shared bPC and decay discPC on the supervised training task of Section 4.1. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
Acc % RMSE
Model MNIST Fashion-MNIST MNIST Fashion-MNIST
arbitrary graph PC $71.47^{\pm 0.27}$ $62.50^{\pm 6.05}$ ${0.1685^{\pm 0.0299}}$ $0.2426^{\pm 0.0077}$
shared-bPC $97.39^{\pm 0.08}$ $87.99^{\pm 0.04}$ $0.1380^{\pm 0.0042}$ $0.2560^{\pm 0.0093}$
decay discPC - - ${0.4047^{\pm 0.0005}}$ $0.3433^{\pm 0.0002}$
Figure 8: MNIST (left) and Fashion-MNIST (right) class average image generation for agPC, shared bPC and decay discPC.
Appendix D Balancing discriminative and generative energies in bPC
In this experiment, we investigate how the relative weighting of the discriminative and generative energy terms ($\alpha_{disc}/\alpha_{gen}$) affects bPC’s learning in the supervised and unsupervised settings described in Sections 4.1 and 4.2. For each setting, we perform hyperparameter tuning of the bPC model while constraining the ratio $\alpha_{disc}/\alpha_{gen}$, and report the test performance of the model with the optimal parameters. We use a grid search over the following hyperparameters: the activity learning rate [0.01, 0.003, 0.001], the parameter learning rate [0.001, 0.0003, 0.0001], and the weight decay [0., 0.0001, 0.0003, 0.001]. Other settings, including a GeLU activation function and an activity momentum of 0, are constant. All remaining parameters, such as the number of epochs, batch size, and number of inference steps during training and evaluation, are kept consistent with the experiments described in Sections 4.1 and 4.2.
In the supervised case, generation RMSE remains stable across ratios of $\alpha_{disc}/\alpha_{gen}$, but classification accuracy declines as the discriminative energy is down-scaled. We suspect this is due to a mismatch in energy magnitudes. A typical discPC model has an energy of around 0.1 for test data samples after training, while genPC has an energy of approximately 50. Thus, the generative energy dominates when $\alpha_{disc}$ is not significantly larger than $\alpha_{gen}$, leading to a poor classification like genPC models. In contrast, bPC’s unsupervised learning performance is relatively robust across a wide range of ratios, but degrades sharply when the discriminative energy is too large. We also observed increased training instability when $\alpha_{disc}\gg\alpha_{gen}$.
These results highlight the importance of appropriately scaling the energy terms based on the task. Hybrid tasks such as those in Section 4.3 require tailored weighting across the models to ensure effective learning. Future work could make the energy scaling a learnable parameter. This change could make bPC learn optimal scalings autonomously.
Figure 9: bPC’s supervised (left) and unsupervised (right) learning performance on MNIST depending on the relative weighting of its discriminative and generative energy term.
Appendix E Effect of max-pooling on generative artifacts
In this experiment, we modify the VGG-5 architecture of the models from Section 4.3 by removing max-pooling layers and increasing the convolutional strides from one to two. We also increase the inference time during training to 48 inference steps. All other aspects of the experiment remain unchanged.
As shown in Figure 10, this change eliminates the checkerboard artefacts in the reconstructed images and improves reconstruction quality for both bPC and hybridPC. This improvement is reflected in the lower RMSE values reported in Table 15. However, removing max-pooling reduces classification accuracy across all models, with bPC experiencing a drop of approximately 25%. Despite this, the overall trends remain consistent: bPC show a substantially better discriminative performance than hybridPC for comparable reconstruction RMSE.
Figure 10: CIFAR-10 image reconstruction from latent representation and labels for models of Section 4.3 without max-pooling.
Table 15: Classification accuracy and image reconstruction RMSE from latent representations for CIFAR-10 models considered in Section 4.3 without max-pooling. Higher accuracy and lower RMSE are better. We report the mean +/- sem over five seeds.
Model Acc % RMSE
bPC $61.67^{\pm 0.29}$ $0.2178^{\pm 0.0047}$
hybridPC $37.36^{\pm 0.18}$ $0.2196^{\pm 0.0010}$
AE $84.58^{\pm 0.11}$ $0.1777^{\pm 0.0005}$
hybridBP $85.79^{\pm 0.26}$ ${0.1095^{\pm 0.0018}}$
Appendix F Sample generation of combined discBP + genBP models
In this experiment, we repeat the combined model image generation procedure from Section 4.4 using discBP and genBP. The two models share an input layer, which is iteratively updated to minimise their energy functions until the energy falls below the 10th percentile of each model’s energy distribution on the test set. We generate samples using the same procedure as described previously for the discPC + genPC model. We compute the FID and Inception scores to quantify the generation quality.
Figure 11 displays the generated samples for different classes. While the images resemble the class averages, they also exhibit noise across pixels and less visual diversity than bPC samples. As shown in Table 16, the backpropagation-based combined model yields higher FID and lower Inception scores, indicating that bPC generates samples more consistent with the MNIST distribution than the backpropagation-based approach.
Figure 11: Samples generated for the combined discBP + genBP model.
Table 16: Fréchet inception distance and inception score for samples generated for the combined discBP + genBP model compared to bPC and discPC + genPC. Results are given for mean +/- sem for three seeds.
Model Inception FID
bPC $\mathbf{6.05^{\pm 0.17}}$ $\mathbf{44.4^{\pm 2.2}}$
discPC + genPC $3.62^{\pm 0.03}$ $140.5^{\pm 2.1}$
discBP + genBP $5.00^{\pm 0.02}$ $99.1^{\pm 3.5}$